CSG 399: Approximation Algorithms

Spring 2004 (Instructor: Rajmohan Rajaraman)

The class meeting on January 7 has been canceled. The first class meeting will be on January 14.

Course information and a tentative schedule are available from the following course web page:

http://www.ccs.neu.edu/home/rraj/Courses/G399/S04/www/